OpenStudy (anonymous):

can someone check my answer??? 0 minus 19 minus 25 minus 25 plus 15 = -54 Sarah owns a small business.There was a loss of $19 on thursday and a loss of 23 on friday.On saturday there was a loss of $11,And on sunday there was a profit of $15.Find the total profit or loss for the four days. A.35 loss B.19 profit C.27 loss D.57 profit

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

\(\large\color{blue}{ \bf 0-19-25+25-15= }\) you can cancel 25s \(\large\color{blue}{ \bf 0-19-15= }\) now finish the calculation \(\large\color{blue}{ \bf -34~. }\)

OpenStudy (solomonzelman):

Now, for your next question. "There was a loss of $19 on Thursday and a loss of 23 on Friday. On Saturday there was a loss of $11" i.e. \(\large\color{blue}{ \bf -19-23-11 }\) add to this one time, when she profited $15 on Sunday... \(\large\color{blue}{ \bf -19-23-11+15 }\) SO.... \(\large\color{blue}{ \bf -19-23-11+15= }\) \(\large\color{blue}{ \bf -42-11+15= }\) \(\large\color{blue}{ \bf -53+15= }\) \(\large\color{blue}{ \bf -38~. }\) none of the options are correct -:(

OpenStudy (solomonzelman):

Check your options and the question again...

OpenStudy (anonymous):

can you retype that stuff but with this becuase this is the correct problem Sarah owns a small business.There was a loss of $16 on thursday and a loss of 9 on friday.On saturday there was a loss of $5,And on sunday there was a profit of $27.Find the total profit or loss for the four days. A.35 loss B.19 profit C.27 loss D.57 profit